Suppose that n represents a positive even integer. What expression represents the next even integer? the next odd integer? I am thinking of three consecutive even integers, whose sum is 204. What are they?
(continued) Find three consecutive odd numbers whose sum is 117. Find two ways to do this.

Answer by richwmiller(17219)

x+x+2+x+4=204
3x+6=204
3x=198
x=66
66 68 70
x+x+2+x+4=117
3x+6=117
3x=111
x=37
37,39 41
117/3=39
39 is the middle term
37 39 41